# A new economic and financial theory of money

### v6

@@ VII Solution of the Hamilton-Jacobi-Bellman equation
To add external forces, that is non-conservative forces (that may be globally conservative), we analytically continue the Hamiltonian and make the canonical transformation $\bar{q}=(q+\text{i}\,p)/\sqrt{2}$ and $\bar{p}=(p+\text{i}\,q)/\sqrt{2}$. The complex analytic Hamiltonian $H(\beta)$ is now given by
@@
and one for $H_{\text{Im}}$ (non-conservative motion generated by $\text{Ad}(H)=\text{i}H_{\text{Im}}=\text{i}\omega\tau=\text{i}\theta$ and parameterized by group parameter $\text{i}E/\omega=\text{i}J$) with equations of motion
@@
When the Hamilton-Jacobi-Bellman equation given in Eq. 50 is solved in this analytically continued extended phase space, the transformed analytic Hamiltonian is given by
@@
The equations of motion for the conservative motion with group parameter $\tau$ are
@@
and the equations of motion for the non-conservative motion with group parameter $\text{i}J$ are
@@
Given this theory, we move on to the practical application of it to control the system. This application will use the concepts of Artificial Intelligence (Hastie et al., 2009; Sugiyama, 2015; Goodfellow et al., 2016), as interpreted by Glinsky (2024b, c). First, construct a dataset by either doing an ensemble of simulations of the system or by observing the system. It will be assumed that the system has a small number of dimensions. Most of the systems of interest present themselves in the domain of collective fields $f(x)$ that are elements of a Hilbert space not $q(\tau)$ with a small number of components. How to de-convolute from the domain of the collective motion, the field $f(x)$, to the domain of the individual, $q(\tau)$, (that is from a Hilbert space to $\mathbb{C}^{n}$) using the Heisenberg Scattering Transformation (HST) and a Principal Components Analysis (PCA) will be discussed at the end of this section. This transformation can be done because the collective acts as one, because of the correlation or synchronization specified by the S-matrix $S_{m}$ given in Eq.  91, which are the derivatives of the analytic Hamiltonian $H(\beta)$ of the individual. It is helpful to apply an external force $F_{\text{ext}}(p,q)$ to the system being simulated or observed to sample phase space more efficiently. A good choice would be a dissipation or a random diffusion which samples phase space well, as the system gradually relaxes to the stable equilibriums. It is also good to apply an external force that is constructed to keep the dynamic trajectory in the vicinity of the unstable local maximums, that is stabilizes the unstable equilibriums. The set of variables that should be recorded are $\tau$, in addition to variables that are related to the state $q(\tau)$ and the co-state $p(\tau)$.
@@
Given the dataset, train a neural network with an architecture that matches the structure of the solution to the HJB equation to estimate: (1) the decoding of the $p$ and $q$ coordinates into the $P(p,q)$ and $Q(p,q)$ coordinates that are the solution to the HJB equation as well as the encoding of $P$ and $Q$ to $p(P,Q)$ and $q(P,Q)$, (2) the value function $W(q,P)$ that generates these canonical transformations and is related to the imaginary part of the analytic Hamiltonian as shown in Eq. 83, (3) the mapping of $P$ to the real part of the analytic Hamiltonian $E_{P}(P)$, (4) the frequency $\omega_{Q}(P)$, (5) the policy $\pi(q,P)$ and (6) the analytic advance of $P$ and $Q$ given in Eqns. 79 and 78. Multi Layer Perceptrons (MLPs) (Goodfellow et al., 2016) are used to approximate some of the functions. The derivative functions are calculated by back propagating the MLPs. This architecture is shown in Fig. 9. It is important that Rectified Linear Units (ReLUs) are used as activation functions in the MLPs because the MLPs are approximating analytic functions which are maximally flat but do have a limited number of singularities where the derivative is discontinuous. MLPs with ReLUs are very good at doing this since they are universal piecewise linear approximators with discontinuities in the derivative.
@@
Figure 9: The neural network architecture to estimate the solution of the Hamilton-Jacobi-Bellman equation. It starts with a decoder from $(p,q)$ to $(P,Q)$ generated by the value function $W(q,P)$ related to the imaginary part of the analytic function $H(\beta)$ by Eq. 83, a mapping to the real part of the analytic function $H(\beta)$ given by $E_{P}(P)$, the frequency $\omega_{Q}(P)$, the policy $\pi(q,P)$, the analytic mapping to advance $(P,Q)$, and a encoder from $(P,Q)$ to $(p,q)$. The universal function approximator is a Multi Layer Perceptron (MLP).
@@
There is a non-trivial detail in this training step. Although one has the inputs ($p_{0}$, $q_{0}$, $d\tau$) and outputs ($p$, $q$), what is $dE$? For a conservative system with no external force being applied $dE=0$, but that is not the case with this dataset. The solution is to use the encoder $E_{P}(p,q)$ to estimate $dE=E_{P}(p,q)-E_{P}(p_{0},q_{0})$, using the target outputs as an input to estimate $E_{P}(p,q)$, as shown in Fig. 10. If this workflow is being used to train a surrogate where the external force is part of the dynamics that is being modeled, a model for the external force $F_{\text{ext}}(\omega_{Q},Q)$ needs to be estimated using an MLP so that $dE=\omega_{Q}\,F_{\text{ext}}(\omega_{Q},Q)\,d\tau$, as shown in Fig. 11. If the force is resistive, diffusive friction $F_{\text{ext}}=-\epsilon_{P}\,\omega_{Q}$, where $\epsilon_{P}\ll J_{0}$. In this case, Rayleigh’s Dissipation Function can be defined
@@
Now to the solution to the general problem of control. First, it might be necessary to change the reward or potential economic energy function from $R_{0}(q)$ to $R(q)$ using $F_{c}(q)$ of Eq. 88. This is improving the geometry by changing the location of $\beta^{*}$. The trick is identifying $R_{0}(q)$. This can be done by learning the solution to the HJB equation $E_{P}(P)$ and $W(q,P)=W_{P}(p,q)$, that is the energy or economic activity and the canonical generating function. The solution can be harvested for $R_{0}(q)$, but more importantly for $\beta^{*}$.
@@
Figure 12: Graph showing the equilibriums where $\omega_{Q}(P^{*})=0$. Plotted is the energy $E(P)$ as a function of $P$. The unstable local maximums are indicated by the x-points, and the stable local minimums (Nash Equilibriums) by the o-points. The energy for the system with no dissipation $\nu=0$ is the thin green line. How this appears in $(p,q)$ phase space can be found in Fig. 5. The heavily damped system $\nu\to\infty$ is the thick black line. The stabilized system (with $F_{sp}(\tau)$ or $F_{sf}(P)$) is the thin blue line.
@@
There are also unstable equilibriums where $\omega^{\prime}_{Q}<0$, that is local maximums. These unstable equilibriums are also known as saddle points or x-points or semi-attractors that first attract the trajectory to them with the system staying near them for a long time since $\omega_{Q}=0$ (a meta stable state). But, eventually the system will start to move away from the semi-attractor, being repelled away from it. The trajectory will eventually approach and be attracted back to the semi-attractor or might approach and be attracted to another semi-attractor (or several other semi-attractors) before being repelled by the last semi-attractor and returning to the first semi-attractor. An example of a set of two semi-attractors could be a bull and a bear market. These saddle points will need to be stabilized by application of an external conservative force, then cooled by a matched non-conservative thermal force, otherwise the trajectory will evolve into the basin of attraction of a stable attractor and eventually relax to that attractor. While small amounts of energy are needed to move from semi-attractor to semi-attractor, a significant amount of energy is needed to move from one attractor to another attractor or semi-attractor. One needs to climb out of the basin of attraction or potential well, that is climb back up to the mountain pass (saddle point) out of the basin before descending into another basin. When these semi-attractors are stabilized, there is still a local minimization of the economic action to efficiently accomplish the task, but the sub-economy is kept at a global sustainable maximum.
It should be noted that if $E_{P}\to\infty$, $\omega_{Q}\to 0$, and $\omega^{\prime}_{Q}\to-\infty$, a conservative stabilizing force can not be constructed. Any attempt at stabilization will be a non-conservative force that will be doing work on the system if it is stabilizing it (that is pumping energy into the system). Another way of looking at it is that it would take an infinite force to stabilize the system.
@@
The stabilization and cooling of the saddle points can be done directly via a feedback force
$\boxed{F_{sf}(P)=-\omega_{sf}(P-P^{*})-\epsilon_{P}\,\omega_{Q}(P)=\Delta(dP/d\tau),}$ (99)
@@
where $\omega_{sf}\gtrsim\omega_{0}$ and $\epsilon_{P}\ll J_{0}$, and $\omega_{0}$ is the dominate spectral frequency or the ground state frequency $\omega_{0}=E_{0}/J_{0}$. This is difficult to do because both $P$ (which is oscillating rapidly about $P^{*}$) and $P^{*}$ must be known or measured. It is much better to apply the ponderomotive equivalent and random walk equivalent
$\boxed{F_{sp}(\tau)=f_{0}\,\text{e}^{\text{i}\,\omega_{sp}\tau}+\omega_{0}\,\varepsilon_{P}=\Delta(dP/d\tau),}$ (100)
@@
where $\omega_{0}\ll\omega_{sp}$, $J_{0}\,\omega_{0}\lesssim f_{0}\ll J_{0}\,\omega_{sp}$ (so that the ponderomotive force is large but the motion is small) and $\varepsilon_{P}$ is a random $\Delta P$ of size $\epsilon_{P}\ll J_{0}$ taken every $2\pi/\omega_{0}$. This $F_{sp}(\tau)$ force is not dependant on $P^{*}$ or $P$, just the time invariant mapping generated by $W(q,P)$ of $p(P,Q)$ and $q(P,Q)$, and the functional transformation iPCA+iHST of $f[p(\tau),q(\tau)](x)$ and $\pi[p(\tau),q(\tau)](x)$ or $f[p+\text{i}q](x)+\text{i}\pi[q+\text{i}p](x)$ as will be discussed later in this section.
@@
What the stabilization force has done is to modify and fortify the topology by adding one x-point and one o-point at the location of the original x-point. It has left the o-point where the original x-point was and moved the two x-points out, sandwiching the o-point. This has turned the mountain pass into a high mountain valley. This is shown in Figs. 12, 13, and 14. Note that the Nash Equilibrium at $P^{*}_{\text{o}0}$ is the point of severe economic depression, the Nash Equilibrium at $P^{*}_{\text{o}}$ is the point of economic recession, and the Nash Equilibrium at $P^{*}_{\text{s}}$ is the point of economic prosperity. It is hard to stimulate an economy from an economic recession to economic prosperity since $E_{\text{s}}>E_{\text{o}}$, but it is very hard to stimulate an economy from a severe economic depression to economic prosperity since $E_{\text{s}}\gg E_{\text{o}0}$. Also note that if the economy is not stimulated enough, so that $E=E_{\text{s}}$, the economy will fall back into the recession or depression.
@@
A very profound structure has been put on the dynamics by the sympletic, that is canonical, structure. This symplectic structure can also be viewed as underlying toroidal topologies $T^{2}$ or cylindrical geometries of extended phase space or $\mathbb{C}$. The underlying analytic function $H(\beta)$ will be specified by two types of singular points in the vector field: (1) o-points that are stable local potential energy (that is reward) minimums, and (2) x-points that are unstable local potential energy (that is reward) maximums. The motion is geodesic motion with a metric of the value or action, that is to say the motion takes the path of minimum action (value or rewards or least economic activity or most efficient way to achieve the objective). This asymmetry in stability induces a direction to time and a irreversibility to the motion. The system will relax to states of minimum, not maximum, total energy (that is action or value). This fact can not be altered. What can be altered is the dynamics (topology) so that the state of maximum global (that is total) energy is a state of minimum global energy. Since the Nash Equilibriums are the states of minimum global energy, the topology must be modified so that the global energy maximums are global energy minimums, that is Nash Equilibriums. The topology is modified and fortified by application of the conservative stabilization force. The control force of Eq. 88 enhances the geometry by changing the location of these new desirable Nash Equilibriums. This theory recognizes that systems fundamentally minimize costs, but one person’s costs are the another person’s rewards. In a game, to maximize personal rewards, the game must be played (that is modified) to maximize all players rewards. If the game is played to greedily minimize personal costs, that is to minimize all player’s rewards, personal rewards will be minimized – the unmodified Nash Equilibrium. This is viewing the system as a zero sum game instead of a “rising tide floats all ships” situation. The same is true of negotiations. The best result is a win-win solution, not a win-lose solution. The win-lose situation is really a lose-lose situation.
From the perspective of the electronic currency management firm, the fiscal (investment and operations) and monetary policy has two parts. The first is changing the economic potential or reward $R(q)$ to include all benefits and costs to society through application of a control force $F_{c}(q)$. The second is stabilizing the economic equilibrium (that is fiscal and monetary policy $\pi^{*}(q)$) and reducing the economic fluctuations about the equilibrium by applying a stabilizing force, $F_{sf}(P)$ or more likely $F_{sp}(\tau)$, through arbitrage trading.
@@
The problem with current attempts to control complex systems is not knowing what reward $R_{0}(q)$ the system is naturally optimizing, and not knowing the equilibrium point $P^{*}$ of the system optimizing $R(q)$ (equivalently the equilibrium value $V^{*}(q)$ or the equilibrium policy $\pi^{*}(q)$). Knowing both $F_{0}(q)$ and $\pi^{*}(q)$ are essential to controlling the system to optimize $R(q)$ and to be stable with minimum fluctuations about the equilibrium where the objective $V(q,P,\tau)$ is optimized. The conservative force that must be applied is $F_{c}(q)$ and the stabilizing and cooling force is $F_{sf}(P)$. A simple way to state this is that the controller needs to know what to control about. In this case, it is $F_{0}(q)$ and $\pi^{*}(q)$ or $P^{*}$. For the ponderomotive control with $F_{sp}(\tau)$, current attempts at control do not know the canonical transformation generated by $W(q,P)$ or the characteristic spectrums $\left|\beta_{i}(z)\right>$ that need to be applied to the fields, as will be discussed next in this section.
@@
Now to the postponed, but important issue, that the system does not present itself in the low finite dimensional dynamical state $q$ and canonically conjugate momentum $p$ or co-state of the individual, but as an infinite dimensional field $f(x)$ and its canonically conjugate field momentum $\pi(x)$ of the collective (Glinsky, 2024b, c), where $x$ is the base manifold or space with metric $\tau(x)$. Another way of looking at this is that nature presents itself as a convoluted form of the dynamical variables that needs to be de-convoluted or decoded into the dynamical variables. It is very common that $x=t$ and $\tau(t)=t$, but could be as sophisticated as $x=(x,y,z,t)$ and $\tau(x,y,z,t)=\sqrt{t^{2}-(x^{2}+y^{2}+z^{2})/c^{2}}$ (that is relativistic dynamics on space-time). The transformation from the infinite dimensional Hilbert space with coordinates $(\pi,f)$ to the finite dimensional $\mathbb{C}^{n}$ space with coordinates $(p,q)$, has been an ongoing challenge to physics and complex system analysis. It has manifested itself as the renormalization challenge to physicists that was first addressed by Ken Wilson (Wilson, 1971) and has been the subject of two Nobel prizes. Paul Dirac felt that renormalization is “not a logical mathematical process” (Dirac, 1982). There have been recent developments on this subject that have approached renormalization as a logical mathematical process, culminating in the Heisenberg Scattering Transformation (HST). The HST is a closed form, specified Convolutional Neural Network (CNN) or Wavelet Conditional Renormalization Group (WC-RG) (Marchand et al., 2022) which is a transformation from canonical $(\pi,f)$-field phase space to $\mathbb{C}^{n}$ with coordinates $\beta$. When combined with the specification of the analytic function $H(\beta)$ or the equilibrium points (topological homology classes) $\beta^{*}$, it is a Generative Pretrained Transformer (GPT) (Li et al., 2022), and when further deployed as a controller it is a Deep Reinforcement Learning (DRL) (Bertsekas and Tsitsiklis, 1996; Sutton and Barto, 2018; Yoon et al., 2021; Wang et al., 2021; Mnih et al., 2015). The HST will be used as a pre-processor, and the inverse HST (iHST) will be used as a post-processor to the previously discussed analysis, as shown in Fig. 16. A detailed discussion of the relationship of our optimal control methodology to DRL can be found in Sec. X, where particular attention is paid to practical differences between our optimal control methodology and DRL.
@@
Details of the HST can be found in Glinsky and Maupin (Glinsky and Maupin, 2023; Glinsky, 2024a, 2023a, 2023b), along with an example of how it can be combined with an MLP and PCA to form a surrogate model for resistive MagnetoHydroDynamics. The equation for the HST is

### v7

@@ VII Solution of the Hamilton-Jacobi-Bellman equation
To add external forces, that do not change the conservative nature of the system, we analytically continue the Hamiltonian and make the canonical transformation $\bar{q}=(q+\text{i}\,p)/\sqrt{2}$ and $\bar{p}=(p+\text{i}\,q)/\sqrt{2}$. The complex analytic Hamiltonian $H(\beta)$ is now given by
@@
and one for $H_{\text{Im}}$ (the motion generated by $\text{Ad}(H)=\text{i}H_{\text{Im}}=\text{i}\omega\tau=\text{i}\theta$ and parameterized by group parameter $\text{i}E/\omega=\text{i}J$) with equations of motion
@@
When the Hamilton-Jacobi-Bellman equation given in Eq. (50) is solved in this analytically continued extended phase space, the transformed analytic Hamiltonian is given by
@@
The equations of motion for the conservative motion, with group parameter $\tau$, are
@@
and the equations of motion due to an external force doing work on the system, with group parameter $\text{i}J$, are
@@
Given this theory, we move on to the practical application of it to control the system. This application will use the concepts of Artificial Intelligence (Hastie et al., 2009; Sugiyama, 2015; Goodfellow et al., 2016), as interpreted by Glinsky (2024d, e). First, construct a dataset by either doing an ensemble of simulations of the system or by observing the system. It will be assumed that the system has a small number of dimensions. Most of the systems of interest present themselves in the domain of collective fields $f(x)$ that are elements of a Hilbert space not $q(\tau)$ with a small number of components. How to de-convolute from the domain of the collective motion, the field $f(x)$, to the domain of the individual, $q(\tau)$, (that is from a Hilbert space to $\mathbb{C}^{n}$) using the Heisenberg Scattering Transformation (HST) and a Principal Components Analysis (PCA) will be discussed at the end of this section. This transformation can be done because the collective acts as one, because of the correlation or synchronization specified by the S-matrix $S_{m}$ given in Eq.  91, which are the derivatives of the analytic Hamiltonian $H(\beta)$ of the individual.
Start by doing an ensemble of simulations or measurements on the system of interest. It is helpful to apply an external force $F_{\text{ext}}$ to the system being simulated or observed to sample phase space more efficiently. A good choice would be a dissipation or a random diffusion which samples phase space well, as the system gradually relaxes to the stable equilibriums. It is also good to apply an external force that is constructed to keep the dynamic trajectory in the vicinity of the unstable local maximums, that is stabilizes the unstable equilibriums. The set of variables that should be recorded are $\tau$, in addition to variables that are related to the state $q(\tau)$ and the co-state $p(\tau)$.
@@
Given the dataset, train a neural network with an architecture that matches the structure of the solution to the HJB equation to estimate: (1) the decoding of the $p$ and $q$ coordinates into the $P(p,q)$ and $Q(p,q)$ coordinates that are the solution to the HJB equation as well as the encoding of $P$ and $Q$ to $p(P,Q)$ and $q(P,Q)$, (2) the value function $W(q,P)$ that generates these canonical transformations and is related to the imaginary part of the analytic Hamiltonian as shown in Eq. (83), (3) the mapping of $P$ to the real part of the analytic Hamiltonian $E_{P}(P)$, (4) the frequency $\omega_{Q}(P)$, (5) the policy $\pi(q,P)$ and (6) the analytic advance of $P$ and $Q$ given in Eqns. (79) and (78). Multi Layer Perceptrons (MLPs) (Goodfellow et al., 2016) are used to approximate some of the functions. The derivative functions are calculated by back propagating the MLPs. This architecture is shown in Fig. 9. It is important that Rectified Linear Units (ReLUs) are used as activation functions in the MLPs because the MLPs are approximating analytic functions which are maximally flat but do have a limited number of singularities where the derivative is discontinuous. MLPs with ReLUs are very good at doing this since they are universal piecewise linear approximators with discontinuities in the derivative.
@@
Figure 9: The neural network architecture to estimate the solution of the Hamilton-Jacobi-Bellman equation. It starts with a decoder from $(p,q)$ to $(P,Q)$ generated by the value function $W(q,P)$ related to the imaginary part of the analytic function $H(\beta)$ by Eq. (83), a mapping to the real part of the analytic function $H(\beta)$ given by $E_{P}(P)$, the frequency $\omega_{Q}(P)$, the policy $\pi(q,P)$, the analytic mapping to advance $(P,Q)$, and a encoder from $(P,Q)$ to $(p,q)$. The universal function approximator is a Multi Layer Perceptron (MLP).
@@
There is a non-trivial detail in this training step. Although one has the inputs ($p_{0}$, $q_{0}$, $d\tau$) and outputs ($p$, $q$), what is $dE$? For a conservative system with no external force being applied $dE=0$, but that is not the case with this dataset. The solution is to use the decoder $E_{P}(p,q)$ to estimate $dE=E_{P}(p,q)-E_{P}(p_{0},q_{0})$, using the target outputs as an input to estimate $E_{P}(p,q)$, as shown in Fig. 10. If this workflow is being used to train a surrogate where the external force is part of the dynamics that is being modeled, a model for the external force $F_{\text{ext}}(\omega_{Q},Q)$ needs to be estimated using an MLP so that $dE=\omega_{Q}\,F_{\text{ext}}(\omega_{Q},Q)\,d\tau$, as shown in Fig. 11. If the force is resistive, diffusive friction $F_{\text{ext}}=-\epsilon_{P}\,\omega_{Q}$, where $\epsilon_{P}\ll J_{0}$. In this case, Rayleigh’s Dissipation Function can be defined
@@
Now to the solution to the general problem of control. First, it might be necessary to change the reward or potential economic energy function from $R_{0}(q)$ to $R(q)$ using $F_{c}(q)$ of Eq. (88). This is improving the geometry by changing the location of $\beta^{*}$. The trick is identifying $R_{0}(q)$. This can be done by learning the solution to the HJB equation $E_{P}(P)$ and $W(q,P)=W_{P}(p,q)$, that is the energy or economic activity and the canonical generating function. The solution can be harvested for $R_{0}(q)$, but more importantly for $\beta^{*}$.
@@
Figure 12: Graph showing the equilibriums where $\omega_{Q}(P^{*})=0$. Plotted is the energy $E(P)$ as a function of $P$. The unstable local maximums are indicated by the x-points, and the stable local minimums (Nash Equilibriums) by the o-points. The energy for the system with no dissipation $\nu=0$ is the thin green line. How this appears in $(p,q)$ phase space can be found in Fig. 5. The heavily damped system $\nu\to\infty$ is the thick black line. The stabilized system (with $F_{\text{sp}}(\tau)$ or $F_{\text{sf}}(P)$) is the thin blue line.
@@
There are also unstable equilibriums where $\omega^{\prime}_{Q}<0$, that is local maximums. These unstable equilibriums are also known as saddle points or x-points or semi-attractors that first attract the trajectory to them with the system staying near them for a long time since $\omega_{Q}=0$ (a meta stable state). But, eventually the system will start to move away from the semi-attractor, being repelled away from it. The trajectory will eventually approach and be attracted back to the semi-attractor or might approach and be attracted to another semi-attractor (or several other semi-attractors) before being repelled by the last semi-attractor and returning to the first semi-attractor. An example of a set of two semi-attractors could be a bull and a bear market. These saddle points will need to be stabilized by application of an external force, then cooled by an external thermal force, otherwise the trajectory will evolve into the basin of attraction of a stable attractor and eventually relax to that attractor. While small amounts of energy are needed to move from semi-attractor to semi-attractor, a significant amount of energy is needed to move from one attractor to another attractor or semi-attractor. One needs to climb out of the basin of attraction or potential well, that is climb back up to the mountain pass (saddle point) out of the basin before descending into another basin. When these semi-attractors are stabilized, there is still a local minimization of the economic action to efficiently accomplish the task, but the sub-economy is kept at a global sustainable maximum.
@@
The stabilization and cooling of the saddle points can be done directly via an external feedback force
$\boxed{F_{\text{sf}}(P)=-\omega_{\text{sf}}(P-P^{*})-\epsilon_{P}\,\omega_{Q}(P)=\Delta(dP/d\tau),}$ (99)
@@
where $\omega_{\text{sf}}\gtrsim\omega_{0}$ and $\epsilon_{P}\ll J_{0}$, and $\omega_{0}$ is the dominate spectral frequency or the ground state frequency $\omega_{0}=E_{0}/J_{0}$. This is difficult to do because both $P$ (which is oscillating rapidly about $P^{*}$) and $P^{*}$ must be known or measured. It is much better to apply the ponderomotive equivalent and random walk equivalent
$\boxed{F_{\text{sp}}(\tau)=f_{0}\,\text{e}^{\text{i}\,\omega_{\text{sp}}\tau}+\omega_{0}\,\varepsilon_{P}=\Delta(dP/d\tau),}$ (100)
@@
where $\omega_{0}\ll\omega_{\text{sp}}$, $J_{0}\,\omega_{0}\lesssim f_{0}\ll J_{0}\,\omega_{\text{sp}}$ (so that the ponderomotive force is large but the motion is small) and $\varepsilon_{P}$ is a random $\Delta P$ of size $\epsilon_{P}\ll J_{0}$ taken every $2\pi/\omega_{0}$. This external $F_{\text{sp}}(\tau)$ force is not dependant on $P^{*}$ or $P$, just the time invariant mapping generated by $W(q,P)$ of $p(P,Q)$ and $q(P,Q)$, and the functional transformation iPCA+iHST of $f[p(\tau),q(\tau)](x)$ and $\pi[p(\tau),q(\tau)](x)$ or $f[p+\text{i}q](x)+\text{i}\pi[q+\text{i}p](x)$ as will be discussed later in this section.
@@
What the stabilization force has done is to modify and fortify the topology by adding one x-point and one o-point at the location of the original x-point. It has left the o-point where the original x-point was and moved the two x-points out, sandwiching the o-point. This has turned the mountain pass into a high mountain valley. This is shown in Figs. 12, 13, and 14. Note that the Nash Equilibrium at $P^{*}_{\text{o}0}$ is the point of severe economic depression, the Nash Equilibrium at $P^{*}_{\text{o}}$ is the point of economic recession, and the unstable equilibrium at $P^{*}_{\text{s}}$ is the point of economic prosperity. It is hard to stimulate an economy from an economic recession to economic prosperity since $E_{\text{s}}>E_{\text{o}}$, but it is very hard to stimulate an economy from a severe economic depression to economic prosperity since $E_{\text{s}}\gg E_{\text{o}0}$. Also note that if the economy is not stimulated enough, so that $E=E_{\text{s}}$, the economy will fall back into the recession or depression.
@@
A very profound structure has been put on the dynamics by the sympletic, that is canonical, structure. This symplectic structure can also be viewed as underlying toroidal topologies $T^{2}$ or cylindrical geometries of extended phase space or $\mathbb{C}$. The underlying analytic function $H(\beta)$ will be specified by two types of singular points in the vector field: (1) o-points that are stable local potential energy (that is reward) minimums, and (2) x-points that are unstable local potential energy (that is reward) maximums. The motion is geodesic motion with a metric of the value or action, that is to say the motion takes the path of minimum action (value or rewards or least economic activity or most efficient way to achieve the objective). This asymmetry in stability induces a direction to time and a irreversibility to the motion. The system will relax to states of minimum, not maximum, total energy (that is action or value). This fact can not be altered. What can be altered is the dynamics (topology) so that the state of maximum global (that is total) energy is a state of minimum global energy. Since the Nash Equilibriums are the states of minimum global energy, the topology must be modified so that the global energy maximums are global energy minimums, that is Nash Equilibriums. The topology is modified and fortified by application of the conservative stabilization force. The control force of Eq. (88) enhances the geometry by changing the location of these new desirable Nash Equilibriums. This theory recognizes that systems fundamentally minimize costs, but one person’s costs are the another person’s rewards. In a game, to maximize personal rewards, the game must be played (that is modified) to maximize all players rewards. If the game is played to greedily minimize personal costs, that is to minimize all player’s rewards, personal rewards will be minimized – the unmodified Nash Equilibrium. This is viewing the system as a zero sum game instead of a “rising tide floats all ships” situation. The same is true of negotiations. The best result is a win-win solution, not a win-lose solution. The win-lose situation is really a lose-lose situation.
From the perspective of the electronic currency management firm, the fiscal (investment and operations) and monetary policy has two parts. The first is changing the economic potential or reward $R(q)$ to include all benefits and costs to society through application of a control force $F_{c}(q)$. The second is stabilizing the economic equilibrium (that is fiscal and monetary policy $\pi^{*}(q)$) and reducing the economic fluctuations about the equilibrium by applying a stabilizing force, $F_{\text{sf}}(P)$ or more likely $F_{\text{sp}}(\tau)$, through arbitrage trading.
@@
The problem with current attempts to control complex systems is not knowing what reward $R_{0}(q)$ the system is naturally optimizing, and not knowing the equilibrium point $P^{*}$ of the system optimizing $R(q)$ (equivalently the equilibrium value $V^{*}(q)$ or the equilibrium policy $\pi^{*}(q)$). Knowing both $F_{0}(q)$ and $\pi^{*}(q)$ are essential to controlling the system to optimize $R(q)$ and to be stable with minimum fluctuations about the equilibrium where the objective $V(q,P,\tau)$ is optimized. The conservative force that must be applied is $F_{c}(q)$ and the external stabilizing and cooling force is $F_{\text{sf}}(P)$. A simple way to state this is that the controller needs to know what to control about. In this case, it is $F_{0}(q)$ and $\pi^{*}(q)$ or $P^{*}$. For the ponderomotive control with $F_{\text{sp}}(\tau)$, current attempts at control do not know the canonical transformation generated by $W(q,P)$ or the characteristic spectrums $\left|\beta_{i}(z)\right>$ that need to be applied to the fields, as will be discussed next in this section.
@@
Now to the postponed, but important issue, that the system does not present itself in the low finite dimensional dynamical state $q$ and canonically conjugate momentum $p$ or co-state of the individual, but as an infinite dimensional field $f(x)$ and its canonically conjugate field momentum $\pi(x)$ of the collective (Glinsky, 2024d, e), where $x$ is the base manifold or space with metric $\tau(x)$. Another way of looking at this is that nature presents itself as a convoluted form of the dynamical variables that needs to be de-convoluted or decoded into the dynamical variables. It is very common that $x=t$ and $\tau(t)=t$, but could be as sophisticated as $x=(x,y,z,t)$ and $\tau(x,y,z,t)=\sqrt{t^{2}-(x^{2}+y^{2}+z^{2})/c^{2}}$ (that is relativistic dynamics on space-time). The transformation from the infinite dimensional Hilbert space with coordinates $(\pi,f)$ to the finite dimensional $\mathbb{C}^{n}$ space with coordinates $(p,q)$, has been an ongoing challenge to physics and complex system analysis. It has manifested itself as the renormalization challenge to physicists that was first addressed by Ken Wilson (Wilson, 1971) and has been the subject of two Nobel prizes. Paul Dirac felt that renormalization is “not a logical mathematical process” (Dirac, 1982). There have been recent developments on this subject that have approached renormalization as a logical mathematical process, culminating in the Heisenberg Scattering Transformation (HST). The HST is a closed form, specified Convolutional Neural Network (CNN) or Wavelet Conditional Renormalization Group (WC-RG) (Marchand et al., 2022) which is a transformation from canonical $(\pi,f)$-field phase space to $\mathbb{C}^{n}$ with coordinates $\beta$. When combined with the specification of the analytic function $H(\beta)$ or the equilibrium points (topological homology classes) $\beta^{*}$, it is a Generative Pretrained Transformer (GPT) (Li et al., 2022), and when further deployed as a controller it is a Deep Reinforcement Learning (DRL) (Bertsekas and Tsitsiklis, 1996; Sutton and Barto, 2018; Yoon et al., 2021; Wang et al., 2021; Mnih et al., 2015). The HST will be used as a pre-processor, and the inverse HST (iHST) will be used as a post-processor to the previously discussed analysis, as shown in Fig. 16. A detailed discussion of the relationship of our optimal control methodology to DRL can be found in Sec. X, where particular attention is paid to practical differences between our optimal control methodology and DRL.
@@
Details of the HST can be found in Glinsky (2024b). The equation for the HST is